### v1

@@
The study of the high-energy (MeV-GeV) part of GRBs spectrum can play a crucial role in investigating the physics of the prompt emission, but it is often hampered by low statistic and the paucity of GeV observations. In this work, we analyze the prompt emission spectra of the 22 brightest GRBs which have been simultaneously observed by Fermi/GBM and Fermi/LAT, spanning 6 orders of magnitude in energy. The high–energy photon spectra can be modelled with a power-law $N(E)\propto E^{-\beta}$ possibly featuring an exponential cutoff. We find that, with the inclusion of the LAT data, the spectral index $\beta$ is softer than what typically inferred from the analysis of Fermi/GBM data alone. Under the assumption that the emission is synchrotron, we derive the index $p\sim 2.79$ of the power-law energy distribution of accelerated particles ($N(\gamma)\propto\gamma^{-p}$). In 9 out of 22 GRB spectra, we find a significant presence of an exponential cut-off at high-energy, ranging between 14 and 298 MeV. By interpreting the observed cut-off as a sign of pair-production opacity, we estimate the jet bulk Lorentz factor $\Gamma$, finding values in the range 130–330. These values are consistent with those inferred from the afterglow light curve onset time. Finally, by combining the information from the high-energy prompt emission spectrum with the afterglow lightcurve, we provide a method to derive the distance $R$ from the central engine where the prompt emission occurs. These results highlight the importance of including high-energy data, when available, in the study of prompt spectra and their role in addressing the current challenges of the GRB standard model.
Insights into the physics of GRBs
@@
M. E. Ravasio Insights into the physics of GRBs from the high-energy extension of their prompt emission spectraInsights into the physics of GRBs from the high-energy extension of their prompt emission spectraInsights into the physics of GRBs from the high-energy extension of their prompt emission spectraInsights into the physics of GRBs from the high-energy extension of their prompt emission spectra    G. Ghirlanda Insights into the physics of GRBs from the high-energy extension of their prompt emission spectraInsights into the physics of GRBs from the high-energy extension of their prompt emission spectra    G. Ghisellini Insights into the physics of GRBs from the high-energy extension of their prompt emission spectraInsights into the physics of GRBs from the high-energy extension of their prompt emission spectra
(Received xxx / Accepted: xxx)
@@ 1 Introduction
Despite over 50 years of observations, the physical mechanism responsible for the prompt emission of Gamma-Ray Bursts (GRBs) remains poorly understood, hindering our understanding of the Universe’s most energetic and enigmatic phenomena. The fast variability ($\sim$ tens/hundreds of ms) of the prompt emission lightcurve, pointing towards a quite compact region, and the high luminosity of $\gamma$-ray photons of the prompt spectrum were the first pieces of observational evidence supporting the relativistic nature of GRBs, as their combination make the compactness (defined as the total luminosity L emitted over the size R of the emitting region) too high to be compatible with observations (e.g. Baring & Harding 1997; Piran 1999; Lithwick & Sari 2001). A relativistic motion, with $\Gamma\gtrsim 100$ for typical GRBs, is needed to overcome the compactness problem and it implies that we observe blue shifted photons which are significantly softer in the co-moving rest frame. The study of the prompt emission spectra offers a powerful tool to investigate the origin of the radiative mechanism, taking into account the relativistic motion. Thanks to the analysis of large samples of GRBs detected by the Burst And Transient Source Experiment (BATSE, $\sim$ 20 keV – 2 MeV) on board the Compton Gamma Ray Observatory (CGRO), the Gamma-ray Burst Monitor (GBM, $\sim$ 8 keV – 40 MeV) on board Fermi  the Burst Alert Telescope (BAT, 15 – 150 keV) and the X–ray Telescope (XRT, $\sim$ 0.2  – 10  keV) on board Swift, the typical observed GRB prompt spectrum appears non-thermal and it is often described by two power-law functions, with indices $\alpha\,\sim\,-1$ and $\beta\,\sim\,-2.5$ at low- and high-energies, respectively, smoothly connected at a characteristic (observer-frame) peak energy, $E_{\rm peak}$$\sim$ 300 – 500 keV, in the $\nu F_{\nu}$ representation (e.g. Preece et al. 2000; Kaneko et al. 2006; Nava et al. 2011; Goldstein et al. 2012; Gruber et al. 2014).
@@
However, the combination of poor statistics of the prompt spectrum at high energies (steep power–law with a typical $\beta\sim-2.5$ of the Band function) and the low effective area of the available detectors makes it difficult to characterize precisely the spectrum above the peak energy. Besides the first few detections of radiation above several tens of MeV provided by the CGRO/EGRET (Schneid et al. 1992, 1995) and AGILE/GRID (Giuliani et al. 2008; Moretti et al. 2009; Longo et al. 2009) instruments, most of our present knowledge on GRB high–energy emission comes from the Large Area Telescope (LAT, 100 MeV – 300 GeV) on board Fermi. Its larger field of view and effective area with respect to its predecessors allow to detect on average 14 GRBs per year, corresponding to $\sim$ 12% of the GBM-detected bursts (see Nava 2018 for a recent review on the high–energy emission of GRBs). Usually, the LAT data reveal the presence of a high energy emission component, which rises with a small delay ($\sim$ a few seconds) after the prompt emission, it lasts longer, and it is less variable. This emission is typically interpreted as afterglow, originating from shocks between the fireball and the interstellar medium (i.e. external shocks, see e.g. Ghirlanda et al. 2010; Ackermann et al. 2010). Instead, the early emission detected by LAT, simultaneous to the prompt phase and characterized by variability, can be interpreted as of internal origin (e.g. Maxham et al. 2011; Zhang et al. 2011), offering the exquisite opportunity to investigate the high–energy extension of the prompt emission spectrum. In particular, the inclusion of the LAT Low Energy (LLE, 30–100 MeV) is crucial in bridging the gap between the GBM and the LAT data (¿100 MeV), allowing a continuous characterization of the prompt emission spectrum.
@@
When combining GBM and LAT data, the high–energy part of the prompt spectrum has been found to lie on the extrapolation of the keV–MeV (GBM) one, thus consisting of a single power–law up to $\sim$ GeV energies, although with a softer index than the one inferred from fitting the GBM data alone (e.g. Fermi Large Area Telescope Team et al. 2012; Abdo et al. 2009b; Axelsson et al. 2012). Assuming that the spectrum of accelerated electrons is a power–law and that the cooling is dominated by the synchrotron emission, the photon index of the power–law describing the spectrum above the peak energy allows to constrain the spectral index $p$ of the electrons’ energy distribution, providing fundamental insights on the acceleration mechanism (Baring 2009; Shen et al. 2006). On the other hand, it was also shown that the spectrum of a handful of GRBs requires a high–energy break between the GBM and LAT energy bands, revealing the presence of an exponential spectral cutoff from a few tens to hundreds of MeV (Fermi Large Area Telescope Team et al. 2012; Tang et al. 2015; Vianello et al. 2018). The exponential drop in the spectrum is an important diagnostic tool as it could be interpreted as a sign of pair-production opacity, allowing to estimate the bulk Lorentz factor $\Gamma$ of the jet (e.g. Lithwick & Sari 2001) and its compactness, or, in alternative, it could be linked to the maximum energy of accelerated particles. The knowledge of bulk Lorentz factor $\Gamma$ is crucial, as it allows to infer the intrinsic properties of the emitting region, shedding light on many interesting - but still unknown - physical quantities (e.g. the location of the emitting region, the ejecta mass, the typical frequencies of the emitted radiation) that are fundamental to discriminate among different theoretical scenarios. However, it is particularly difficult to obtain valuable constraints on $\Gamma$ from observations. So far, most of the measurements and lower limits currently available on $\Gamma$ are derived from the detection of (or upper limit on) the peak time of the afterglow emission (see e.g. Ghirlanda et al. 2018). The detection of an exponential cut-off in the prompt emission spectrum provides instead an alternative method to derive the bulk Lorentz factor $\Gamma$ of the jet. Moreover, the compactness argument is independent from the radiation mechanism producing the observed gamma-ray emission, giving constraints on $\Gamma$ and related emission parameters which have to be explained by any proposed prompt emission mechanism.
@@ 2 The sample
Fig. 1 shows the fluence measured in the 10-1000 keV energy range (i.e. detected by GBM) vs. the significance of the signal in the range 30-100 MeV (i.e. LLE data), in units of $\sigma$, for the 91 GRBs of the LAT Catalog detected in the LLE energy range, represented by the red points. The selected 22 GRBs are shown by the blue points. The selection criteria described above are represented by the two black dashed lines. From our selection we excluded GRB 090926A (represented by the red point falling in the selected region in Fig. 1). This GRB has an additional high-energy power-law component detected by the LAT during the prompt phase, which also shows a spectral cutoff at 1.4 GeV (Ackermann et al. 2011; Yassine et al. 2017). This prevents the identification of a possible high–energy cutoff on the main spectral component and the accurate determination of the slope of the high–energy power–law, which are both main goals of this work.
Figure 1: The sample of 22 GRBs analyzed (blue points) among the 91 events (red points) detected in the LLE energy range (30–100  MeV), in 10 years of Fermi activity, according to the LAT Catalog. We selected the events applying a cut on the GBM fluence in the 10-1000 keV energy range and a cut on the significance of the signal in the LLE data, both represented by the two black dashed lines. The red point falling in the selected region represents GRB 090926A, which has been excluded from the selection (see the text for an explanation).
@@
Table 1: The sample of 22 long GRBs selected for this work. The events satisfy two selection criteria: a significance in the LLE [30–100 MeV] signal $\sigma\geq$ 20, and a fluence [10–1000 keV] over the total duration of the burst $F>10^{-5}\,\rm erg\,cm^{-2}$, as reported in the LAT and in the GBM Catalog, respectively. The first column lists the name of the GRB (with the last three digits referring to the naming convention of Fermi GBM), the second column reports the fluence of the burst, the third column reports the significance of the signal in the LLE data, in terms of $\sigma$. The last column reports the redshift of the burst, if available. The events are listed in order of decreasing significance of the signal in the LLE data.
[…]
@@ 3 Methods
The GBM is an instrument composed of 12 sodium iodide (NaI, 8 keV–1 MeV) and two bismuth germanate (BGO, 200 keV to 40 MeV) scintillation detectors (Meegan et al. 2009), on board the Fermi satellite. For each GRB, we analysed the data from the two NaI detectors and the BGO detector with the highest count rates. Spectral data files and the corresponding most updated response matrix files (rsp2) are obtained from the online archive  [^3] https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html. We select the energy channels in the range 10–900 keV for NaI detectors, and 0.3–10 MeV for BGO detectors, and exclude the channels in the range 30–40 keV due to the presence of the Iodine K-edge at 33.17 keV [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/GBM_caveats.html. For our sample, which consists of long GRBs, we made use of CSPEC data, which have 1024 ms time resolution. Spectra were extracted with the public software gtburst [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/gtburst.html.
The Large Area Telescope (LAT, Atwood et al. 2009) is a pair-conversion telescope covering the energy range from 30 MeV to more than 300 GeV. The extraction and analysis of LAT data were performed with gtburst, following the procedure described in the online official Fermi guide  [^6] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/gtburst.html and performing an unbinned likelihood analysis. We selected P8R3_TRANSIENT020 class events, filtering photons in the 100 MeV – 10 GeV energy range from a region of interest (ROI) of ${}^{\circ}$ radius centered on the burst location. We applied a maximum zenith angle cut of ${}^{\circ}$ to reduce contamination of gamma-rays from the Earth limb. The LAT Low Energy (LLE) data were retrieved from the Fermi LLE Catalog [^7] https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermille.html and reduced with a similar procedure as for the GBM data through gtburst. The LLE spectra analysed cover the energy range 30-100 MeV.
For each GRB, we performed a joint spectral analysis of the GBM (10 keV – 10 MeV), LLE (30 – 100 MeV) and LAT (100 MeV – 10 GeV) data (see shaded stripes in Fig.2). Spectral analysis is performed with the public software xspec (v. 12.10.1f). For the joint analysis of the GBM, LLE and and LAT data, we used inter–calibration factors among the detectors, scaled to the brightest NaI and free to vary within 30%. We used the PG-Statistic, valid for Poisson data with a Gaussian background, in the fitting procedure.

### v2

@@
Insights into the physics of Gamma-Ray Bursts
@@
M. E. Ravasio Insights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectraInsights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectraInsights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectraInsights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectra    G. Ghirlanda Insights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectraInsights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectra    G. Ghisellini Insights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectraInsights into the physics of Gamma-Ray Bursts from the high-energy extension of their prompt emission spectra
(Received 04/12/2023 / Accepted: 23/02/2024)
The study of the high-energy part (MeV-GeV) of the spectrum of gamma-ray bursts (GRBs) can play a crucial role in investigating the physics of prompt emission, but it is often hampered by low statistics and the paucity of GeV observations. In this work, we analyze the prompt emission spectra of the 22 brightest GRBs which have been simultaneously observed by Fermi/GBM and Fermi/LAT, spanning six orders of magnitude in energy. The high–energy photon spectra can be modeled with a power-law $N(E)\propto E^{-\beta}$ possibly featuring an exponential cutoff. We find that, with the inclusion of the LAT data, the spectral index $\beta$ is softer than what is typically inferred from the analysis of Fermi/GBM data alone. Under the assumption that the emission is synchrotron, we derived a median value of the index $p\sim 2.79$ of the power-law energy distribution of accelerated particles ($N(\gamma)\propto\gamma^{-p}$). In nine out of 22 GRB spectra, we find a significant presence of an exponential cutoff at high energy, ranging between 14 and 298 MeV. By interpreting the observed cutoff as a sign of pair-production opacity, we estimate the jet bulk Lorentz factor $\Gamma$, finding values in the range 130–330. These values are consistent with those inferred from the afterglow light curve onset time. Finally, by combining the information from the high-energy prompt emission spectrum with the afterglow light curve, we exploited a promising method to derive the distance $R$ from the central engine where the prompt emission occurs. The distances ($R>10^{13-15}$ cm) inferred for the only two GRBs in our sample that are suitable for the application of this method, which have only lower limits on their cutoff energies, suggest large emitting regions, although they are still compatible with the standard model. Larger samples of GRBs with measured cutoff energies and afterglow deceleration time will allow for more informative values to be derived. These results highlight the importance of including high-energy data, when available, in the study of prompt spectra and their role in addressing the current challenges of the GRB standard model.
@@ 1 Introduction
Despite over 50 years of observations, the physical mechanism responsible for the prompt emission of gamma-ray bursts (GRBs) remains poorly understood, hindering our understanding of the Universe’s most energetic and enigmatic phenomena. The fast variability (approximately tens or hundreds of milliseconds) of the prompt emission light curve, pointing toward a rather compact region, and the high luminosity of $\gamma$-ray photons of the prompt spectrum were the first pieces of observational evidence supporting the relativistic nature of GRBs, as their combination makes the compactness (defined as the total luminosity L emitted over the size R of the emitting region) too high to be compatible with observations (e.g., Baring & Harding 1997; Piran 1999; Lithwick & Sari 2001). A relativistic motion, with $\Gamma\gtrsim 100$ for typical GRBs, is needed to overcome the compactness problem and it implies that we observed blue-shifted photons that are significantly softer in the co-moving rest frame. The study of the prompt emission spectra offers a powerful tool to investigate the origin of the radiative mechanism, taking the relativistic motion into account. Thanks to the analysis of large samples of GRBs detected by the Burst And Transient Source Experiment (BATSE, $\sim$ 20 keV – 2 MeV) on board the Compton Gamma Ray Observatory (CGRO), the Gamma-ray Burst Monitor (GBM, $\sim$ 8 keV – 40 MeV) on board Fermi , the Burst Alert Telescope (BAT, 15 – 150 keV), and the X–ray Telescope (XRT, $\sim$ 0.2  – 10  keV) on board Swift, the typical observed GRB prompt spectrum appears nonthermal and it is often described by two power-law functions, with indices $\alpha\,\sim\,-1$ and $\beta\,\sim\,-2.5$ at low and high energies, respectively, and they are smoothly connected at a characteristic (observer-frame) peak energy, $E_{\rm peak}$$\sim$ 300 – 500 keV, in the $\nu F_{\nu}$ representation (e.g., Preece et al. 2000; Kaneko et al. 2006; Nava et al. 2011; Goldstein et al. 2012; Gruber et al. 2014).
@@
However, the combination of poor statistics of the prompt spectrum at high energies (steep power law with a typical $\beta\sim-2.5$ of the Band function) and the low effective area of the available detectors makes it difficult to characterize precisely the spectrum above the peak energy. Besides the first few detections of radiation above several tens of MeV provided by the CGRO/EGRET (Schneid et al. 1992, 1995) and AGILE/GRID (Giuliani et al. 2008; Moretti et al. 2009; Longo et al. 2009) instruments, most of our present knowledge on GRB high–energy emission comes from the Large Area Telescope (LAT, 100 MeV – 300 GeV) on board Fermi. Its larger field of view and effective area with respect to its predecessors allow for 14 GRBs to be detected per year on average, corresponding to $\sim$ 12% of the GBM-detected bursts (see Nava 2018 for a recent review on the high–energy emission of GRBs). Usually, the LAT data reveal the presence of a high-energy emission component, which rises with a small delay (approximately a few seconds) after the prompt emission, it lasts longer, and it is less variable. This emission is typically interpreted as afterglow, originating from shocks between the fireball and the interstellar medium (i.e., external shocks, see e.g., Ghirlanda et al. 2010; Ackermann et al. 2010). Instead, the early emission detected by LAT, which is simultaneous with the prompt phase and characterized by variability, can be interpreted as of an internal origin (e.g., Maxham et al. 2011; Zhang et al. 2011), offering the exquisite opportunity to investigate the high–energy extension of the prompt emission spectrum. In particular, the inclusion of the LAT Low Energy (LLE, 30–100 MeV) is crucial in bridging the gap between the GBM and the LAT data (¿100 MeV), allowing a continuous characterization of the prompt emission spectrum.
@@
When combining GBM and LAT data, the high–energy part of the prompt spectrum has been found to lie on the extrapolation of the keV–MeV (GBM) one, thus consisting of a single power law up to approximately GeV energies, although with a softer index than the one inferred from fitting the GBM data alone (e.g., Fermi Large Area Telescope Team et al. 2012; Abdo et al. 2009b; Axelsson et al. 2012). Assuming that the spectrum of accelerated electrons is a power law and that the cooling is dominated by the synchrotron emission, the photon index of the power law describing the spectrum above the peak energy allows one to constrain the spectral index $p$ of the electrons’ energy distribution, providing fundamental insights into the acceleration mechanism (Baring 2009; Shen et al. 2006). On the other hand, it was also shown that the spectrum of a select number of GRBs requires a high–energy break between the GBM and LAT energy bands, revealing the presence of an exponential spectral cutoff from a few tens to hundreds of MeV (Fermi Large Area Telescope Team et al. 2012; Tang et al. 2015; Vianello et al. 2018). The exponential drop in the spectrum is an important diagnostic tool as it could be interpreted as a sign of pair-production opacity, allowing one to estimate the bulk Lorentz factor $\Gamma$ of the jet (e.g., Lithwick & Sari 2001) and its compactness, or, alternatively, it could be linked to the maximum energy of accelerated particles. The knowledge of bulk Lorentz factor $\Gamma$ is crucial, as it allows one to infer the intrinsic properties of the emitting region, shedding light on many interesting - but still unknown - physical quantities (e.g., the location of the emitting region, the ejecta mass, the typical frequencies of the emitted radiation) that are fundamental to discriminate among different theoretical scenarios. However, it is particularly difficult to obtain valuable constraints on $\Gamma$ from observations. So far, most of the measurements and lower limits currently available on $\Gamma$ have been derived from the detection of (or upper limit on) the peak time of the afterglow emission (see e.g., Ghirlanda et al. 2018). The detection of an exponential cutoff in the prompt emission spectrum instead provides an alternative method to derive the bulk Lorentz factor $\Gamma$ of the jet. Moreover, the compactness argument is independent of the radiation mechanism producing the observed gamma-ray emission, giving constraints on $\Gamma$ and related emission parameters that have to be explained by any proposed prompt emission mechanism.
@@ 2 The sample
Fig. 1 shows the fluence measured in the 10-1000 keV energy range (i.e., detected by GBM) versus the significance of the signal in the range 30-100 MeV (i.e., LLE data), in units of $\sigma$, for the 91 GRBs of the LAT Catalog detected in the LLE energy range, represented by the red points. The blue points show the selected 22 GRBs. The selection criteria described above are represented by the two black dashed lines. From our selection, we excluded GRB 090926A (represented by the red point falling in the selected region in Fig. 1). Conversely from the others, this GRB has an additional high-energy power-law component detected by the LAT during the prompt phase, which also shows a spectral cutoff at 1.4 GeV (Ackermann et al. 2011; Yassine et al. 2017). This prevents the identification of a possible high–energy cutoff on the main spectral component and the accurate determination of the slope of the high–energy power–law, which are both main goals of this work.
Figure 1: Sample of 22 GRBs analyzed (blue points) among the 91 events (red points) detected in the LLE energy range (30–100  MeV), in 10 years of Fermi activity, according to the LAT catalog. We selected the events applying a cut on the GBM fluence in the 10-1000 keV energy range and a cut on the significance of the signal in the LLE data, both represented by the two black dashed lines. The red point falling in the selected region represents GRB 090926A, which has been excluded from the selection (see the text for an explanation).
@@
Table 1: Sample of 22 long GRBs selected for this work. The events satisfy two selection criteria: a significance in the LLE [30–100 MeV] signal $\sigma\geq$ 20, and a fluence [10–1000 keV] over the total duration of the burst $F>10^{-5}\,\rm erg\,cm^{-2}$, as reported in the LAT and in the GBM catalog, respectively. The first column lists the name of the GRB (with the last three digits referring to the naming convention of Fermi GBM), the second column reports the fluence of the burst, the third column reports the significance of the signal in the LLE data, in terms of $\sigma$. The last column reports the redshift of the burst, if available. The events are listed in order of decreasing significance of the signal in the LLE data.
GRB Name GBM Fluence Significance of LLE signal Redshift [$10^{-4}$ erg cm-2] $\sigma$ 110721[200] 0.370 61 160625[945] 6.43 49 1.406 080916[009] 0.603 46 4.35 170214[649] 1.77 46 2.53 100724[029] 2.17 40 140206[275] 1.23 39 131108[862] 0.357 37 2.4 141028[455] 0.348 31 2.332 100116[897] 0.334 31 160821[857] 5.22 31 130504[978] 1.29 29 110328[520] 0.192 29 160509[374] 1.79 29 1.17 180720[598] 2.99 24 0.654 151006[413] 0.122 24 160905[471] 0.732 22 150902[733] 0.832 21 100826[957] 1.64 21 090328[401] 0.420 20 0.736 110731[465] 0.229 20 2.83 160910[722] 0.797 20 150202[999] 0.325 20
@@ 3 Methods
For each GRB in our sample, we performed a joint spectral analysis of the GBM (10 keV – 10 MeV), LLE (30 – 100 MeV) and LAT (100 MeV – 10 GeV) data (see shaded stripes in Fig.2). In the following, we describe the data analysis procedure and the fitting models.
3.1 Data analysis
The GBM is an instrument composed of 12 sodium iodide (NaI, 8 keV–1 MeV) and two bismuth germanate (BGO, 200 keV to 40 MeV) scintillation detectors (Meegan et al. 2009), on board the Fermi satellite. For each GRB, we analyzed the data from the two NaI detectors and the BGO detector with the highest count rates. Spectral data files and the corresponding most updated response matrix files (rsp2) are obtained from the online archive  [^3] https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html. We select the energy channels in the range 10–900 keV for NaI detectors, and 0.3–10 MeV for BGO detectors, and exclude the channels in the range 30–40 keV due to the presence of the Iodine K-edge at 33.17 keV [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/GBM_caveats.html. For our sample, which consists of long GRBs, we made use of CSPEC data, which have 1024 ms time resolution. Spectra were extracted with the public software gtburst [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/gtburst.html.
The Large Area Telescope (LAT, Atwood et al. 2009) is a pair-conversion telescope covering the energy range from 30 MeV to more than 300 GeV. The extraction and analysis of LAT data were performed with gtburst, following the procedure described in the online official Fermi guide  [^6] https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/gtburst.html and performing an unbinned likelihood analysis. We selected P8R3_TRANSIENT020 class events, filtering photons in the 100 MeV – 10 GeV energy range from a region of interest (ROI) of 12 radius centered on the burst location. We applied a maximum zenith angle cut of 100 to reduce contamination of gamma-rays from the Earth limb. The LAT Low Energy (LLE) data were retrieved from the Fermi LLE catalog [^7] https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermille.html and reduced with a similar procedure as for the GBM data through gtburst. The LLE spectra analyzed cover the energy range 30-100 MeV.
Spectral analysis is performed with the public software xspec (v. 12.10.1f). For the joint analysis of the GBM, LLE and LAT data, we used inter–calibration factors among the detectors, scaled to the brightest NaI and free to vary within 30%. We used the PG-Statistic, valid for Poisson data with a Gaussian background, in the fitting procedure.